OpenStudy (anonymous):

MEDALS WILL BE GIVING OUT Sam invests $9,400 in a fund yielding 5.2% simple interest annually. At the same time, Pat shops around for a better interest rate and invests $9,400 in a fund yielding 6.0% simple interest annually. After seven years, how much more interest has Pat’s fund earned than Sam’s fund? a. $817.80 b. $3,948.00 c. $652.40 d. $526.40

OpenStudy (anonymous):

@Algorithmic

OpenStudy (princeharryyy):

for sam total si = 7*9400*0.052 = $3421.6 for pat total si = 7*9400*0.06 = $3948 difference = $3948-$3421.6 = $526.4 answer

OpenStudy (princeharryyy):

@DiornJazz

OpenStudy (anonymous):

Thank you @princeharryyy

OpenStudy (princeharryyy):

u r welcome! but, I'm sorry I can't help you with more. I gotta go to sleep. It's 4:0 am. Bye. @DiornJazz

OpenStudy (anonymous):

Bye i wish u could help. Could you inbox me when u awake so i can get more help from u

OpenStudy (princeharryyy):

7 hrs later. Give ma a call @DiornJazz